What is the perimeter of a field tha measures 62 2/3 yards by 48 1/6 yards?

Answer by Boreal(15235)

Add the two
62+48=110
2/3+1/6=5/6
110 5/6 yards
Perimeter is twice the length and width, so it is 221 2/3 yards, multiplying 110 5/6 by 2.